### v2

# An Information-Theoretic Regularizer for Lossy Neural Image Compression

Abstract
Lossy image compression networks aim to minimize the latent entropy of images while adhering to specific distortion constraints. However, optimizing the neural network can be challenging due to its nature of learning quantized latent representations. In this paper, our key finding is that minimizing the latent entropy is, to some extent, equivalent to maximizing the conditional source entropy, an insight that is deeply rooted in information-theoretic equalities. Building on this insight, we propose a novel structural regularization method for the neural image compression task by incorporating the negative conditional source entropy into the training objective, such that both the optimization efficacy and the model’s generalization ability can be promoted. The proposed information-theoretic regularizer is interpretable, plug-and-play, and imposes no inference overheads. Extensive experiments demonstrate its superiority in regularizing the models and further squeezing bits from the latent representation across various compression structures and unseen domains.
1 Introduction
Lossy image compression, or more broadly, lossy data compression, is grounded in a fundamental theorem in information theory: the rate-distortion theorem [1]. Essentially, this theorem establishes a theoretical lower bound for lossy compression and provides valuable guidance for designing practical compression algorithms [2, 3].
Figure 1: (a) Direct coding model; (b) Information diagram for the direct coding, wherein $H(\bm{X})=H(\bm{X}|\bm{\hat{X}})+H(\bm{U})$ and $H(\bm{X})$ is fixed for any known source; (c) Transform coding model (with possible side information $\bm{\hat{Z}}$ [4]).
A basic lossy image compression model is illustrated in Fig.1(a). Let $\bm{X}$, $\bm{\hat{X}}$, and $\bm{U}$ denote the source, the reconstruction, and the discrete index (or the latent), respectively. The encoder is a deterministic mapping $\mathrm{Q}$: $\mathcal{X}^{N}\rightarrow\mathcal{U}$, where $\mathcal{X}^{N}$ denotes a $N$-dimensional image space and $\mathcal{U}=\{1,2,...,M\}$ denotes the discrete index space. The decoder is a bijection mapping $\mathrm{Q}^{-1}$: $\mathcal{U}\rightarrow\mathcal{\hat{X}}^{N}$, where the reconstruction space $\mathcal{\hat{X}}^{N}=\{\bm{c}_{1},...,\bm{c}_{M}\}$ forms a $N$-dimensional codebook. For index $\bm{U}$, entropy coding is applied to losslessly convert the data into the bitstream, which can then be decoded back to $\bm{U}$ on the decoder side.
We may refer to such a model as the direct coding model for its structural simplicity. Subsequently, following the rate-distortion theorem [1], given a bounded distortion measure $d(\cdot)$, and a distortion constraint $D$, the rate of transmitting $\bm{U}$ is lower bounded by the informational rate-distortion function:
$R(D)=\mathop{\text{min}}\limits_{p_{\bm{\hat{X}}|\bm{X}}:\mathbb{E}_{p_{\bm{X}}p_{\bm{\hat{X}}|\bm{X}}}{[d(\bm{X},\bm{\hat{X}})]}<D}\ {\frac{1}{N}I(\bm{X};\bm{\hat{X}})},$ (1)
i.e., the minimum mutual information obtained by traversing all the possible encoder-decoder (or quantizer) designs $p_{\bm{\hat{X}}|\bm{X}}$. Theoretically, this bound can be achieved asymptotically using the vector quantization with $N\rightarrow\infty$ [5]. However, as the encoding complexity grows exponentially with the image dimension $N$, a suboptimal yet more practical solution is the transform coding [6]. As depicted in Fig.1(c), for the transform coding model, a deterministic analysis transform $\mathrm{T}_{\mathrm{A}}$: $\mathcal{X}^{N}\rightarrow\mathcal{Y}^{M}$ and a deterministic synthesis transform $\mathrm{T}_{\mathrm{S}}$: $\mathcal{\hat{Y}}^{M}\rightarrow\mathcal{\hat{X}}^{N}$ are inserted into the compression chain. The analysis transform is designed to decorrelate the source $\bm{X}$, such that a much simpler quantization scheme, e.g., uniform scalar quantization: $\bm{u}_{i}=\mathrm{round}(\bm{y}_{i})$, can be independently applied to each dimension of the transformed coefficients $\bm{Y}$. Building on this framework and leveraging the superior nonlinear capabilities of neural networks, numerous neural image compression networks [7, 4, 8, 9, 10, 11, 12, 13] are proposed, continually pushing the frontier of lossy image compression.
Typically, by parameterizing the $\mathrm{T}_{\mathrm{A}}$ and $\mathrm{T}_{\mathrm{S}}$ in Fig.1(c) as neural networks, we have the following rate-distortion training objective [7]:
$\mathop{\text{min}}\ \underbrace{{\mathbb{E}_{\bm{X}}[-\text{log}\ q_{\phi}(\bm{U})]}}_{R\approx H(\bm{U})}+\lambda\underbrace{\mathbb{E}_{\bm{X}}||\bm{X}-\bm{\hat{X}}||_{2}^{2},}_{D}$ (2)
where the rate $R$ is the estimation of latent entropy $H(\bm{U})$; $q_{\phi}$ is the neural entropy model (or estimated distribution) for $\bm{U}$, whose modeling accuracy can be significantly enhanced by transmitting a small amount of side information $\bm{\hat{Z}}$ (as depicted in Fig.1(c)) as part of the parameters $\phi$ [4]; the distortion $D$ is the $l_{2}$ distance between $\bm{X}$ and $\bm{\hat{X}}$; $\lambda$ is the Lagrangian multiplier, controlling the trade-off of $R$ and $D$. Herein, due to the existence of quantization $\mathrm{Q}(\cdot)$, $\bm{U}$ is discrete in essence. Optimizing the network is then challenging as the gradient is almost zero everywhere for $\mathrm{Q}(\cdot)$. To effectively train the networks, a surrogate is introduced during the training, e.g., additive uniform noise (AUN) [7]:
$\bm{U}=\mathrm{Q}(\bm{Y})\approx\bm{Y}+[-0.5,0.5]^{M},$ (3)
such that the backward gradient of $\mathrm{Q}(\cdot)$ is one. However, as such approximation inevitably results in the train-test mismatch, and thus high gradient estimation error [14], both the efficiency and efficacy of the optimization are hindered. To address this, many research efforts are devoted to designing advanced surrogates and optimization methods [15, 16, 17, 18, 14].
In this paper, we present a completely distinct view on facilitating the optimization of the compression network: regularization. Conceivably, imposing structural constraints on the network optimization can effectively regularize the gradient estimation error, promote convergence, and reduce overfitting. In the context of generative variational autoencoders [19], structural regularization techniques [20, 21, 22, 23, 24, 25] have been effectively applied to improve authentic data generation. However, since neural compression networks largely differ from generative autoencoders in terms of the minimum latent entropy constraint, to the best of our knowledge, no structural regularization methods tailored to this constraint have been explored. In light of this, in this paper, we propose an information-theoretic regularizer for the neural image compression task. In particular, through an in-depth information-theoretic analysis of compression models, we reveal that minimizing the latent entropy $H(\bm{U})$ is, to some extent, equally important as maximizing the conditional source entropy $H(\bm{X}|\bm{\hat{X}})$. Subsequently, by enforcing the network to simultaneously minimize $-H(\bm{X}|\bm{\hat{X}})$, both the in-domain compact representation and out-of-domain generalization ability are experimentally improved.
2 Related works
Neural image compression networks.
Neural image compression networks are primarily composed of transform, quantization, and latent entropy models. The transform module maps the input source into a latent representation so that the dependencies of the source domain can be more easily handled. Unlike the linear transform [26] used in traditional hybrid coding, the neural transform is inherently nonlinear, allowing for more effective decorrelation of the source data. In the literature, various neural transform architectures are proposed. Representative architectures include the recurrent neural network (RNN) [27], convolutional neural network (CNN) [7], non-local attention [28], transformer [29], mixed transformer-CNN [30], state-space models [13] and invertible networks [31]. Subsequently, the continuous data is quantized into discrete representation, wherein the uniform scalar quantization [7] is commonly adopted. However, as vector quantization in principle can better approach the rate-distortion bound (Eqn.(1)), low-complexity vector quantization [15, 32, 33] and dependent scalar quantization [34, 35] schemes are also studied. Finally, once the quantized latent representations are obtained, a neural entropy model is essential for performing entropy coding. Herein, since the latent can not be perfectly decorrelated, an accurate entropy model that can better capture the dependencies and closely match the ground-truth latent distribution is crucial for the overall network performance. Prominent research directions such as spatial-channel contexts mining [8, 9, 36, 12], expressive latent distribution modeling [10, 37, 38] and latent architecture designs [4, 11, 39] are explored.
Optimization over quantized latent representations.
The surrogate-based methods [7, 15, 16, 17, 14] prove to be effective in tackling the non-differentiable quantization function during end-to-end optimization. In addition to AUN [7], various quantization surrogates such as soft assignment [15], Gumbel-softmax trick [16], and soft rounding function [17] have been studied. However, as the surrogate introduces a train-test mismatch in forward calculations, the annealing techniques [15, 16, 17, 14] are then proposed to control the “softness” of surrogates, gradually approximating the hard quantization function as training progresses. The straight-through estimator (STE) [40] provides an alternative approach to address the train-test mismatch by enforcing an identity mapping during the backward calculation, thereby enabling direct hard quantization in the forward pass. Nevertheless, STE ultimately introduces a significant gradient estimation bias for the analysis transform. To counter this, an additional fine-tuning stage for the entropy model and synthesis transform has been suggested [18].
While numerous network optimization methods have been developed from a quantization perspective, the regularization view has largely escaped research attention. To the best of our knowledge, the most related efforts are regularization methods for variational autoencoders (VAEs), such as regularization on latent priors [20, 21, 22], reconstruction [23, 24], and consistency [25]. However, as VAEs are fundamentally designed for generative modeling, these methods are not applicable to optimizing compression networks. To bridge this knowledge gap, this paper proposes a structural regularizer for neural image compression and conducts extensive experiments to demonstrate its superiority.
3 Methodology
In this section, we begin by deriving the equivalent maximization form of the latent entropy minimization problem for both the direct and transform coding models. Subsequently, building on this derivation, we propose a novel regularization method for optimizing the neural compression networks. Note that during our derivations, we use the naive transform coding model without the side information branch (i.e., the $\bm{\hat{Z}}$ depicted in Fig 1(c)). This significantly simplifies our analysis without compromising much practicality, as it has been shown that the information flowing into the side branch is negligible compared to the latent branch [4].
3.1 Motivation: from minimization to maximization
Lemma 1.
Considering a deterministic quantization process ${Q}(\cdot)$ and a deterministic dequantization process ${Q}^{-1}(\cdot)$ for the direct coding model illustrated in Fig. 1(a), the following equalities hold automatically:
$I(\bm{X};\bm{\hat{X}})=H(\bm{\hat{X}})$ (4)
$H(\bm{\hat{X}})=H(\bm{U})$ (5)
$H(\bm{U})=I(\bm{X};\bm{\hat{X}})$ (6)
Proof.
Eqn.(4) follows the facts that $H(\bm{\hat{X}}|\bm{X})=0$, meaning there is no uncertainty in $\bm{\hat{X}}$ given the observation of $\bm{X}$, and $I(\bm{X};\bm{\hat{X}})=H(\bm{\hat{X}})-H(\bm{\hat{X}}|\bm{X})$. Similarly, Eqn.(5) follows the fact that $H(\bm{U}|\bm{\hat{X}})=H(\bm{\hat{X}}|\bm{U})=0$ and $H(\bm{\hat{X}})+H(\bm{U}|\bm{\hat{X}})=H(\bm{U})+H(\bm{\hat{X}}|\bm{U})$. Eqn.(6) is immediately apparent from Eqn.(4) and Eqn.(5). For more details, readers may refer to the Supplementary Material (Sec. 6). ∎
Theorem 1.
For the direct coding model elaborated in Lemma 1, we have
$H(\bm{U})=H(\bm{X})-H(\bm{X}|\bm{\hat{X}}).$ (7)
The informational diagram is thus illustrated as Fig.1(b).
Proof.
This is immediately apparent from Eqn.(6), as the right-hand-side of Eqn.(7) is by definition $I(\bm{X};\bm{\hat{X}})$. ∎
From Theorem 1, for the direct coding model, it is clear that minimizing the latent entropy $H(\bm{U})$ can be equivalently achieved by maximizing the conditional source entropy $H(\bm{X}|\bm{\hat{X}})$, for any known source $\bm{X}$. In practice, this aligns with the training setting of the neural compression network, wherein the training set is given in advance, and thus $H(\bm{X})$ is fixed. However, as the prevailing neural compression networks are based on the transform coding model, in the following discussion, we extend this conclusion to the transform coding setting.
Lemma 2.
Considering a transform coding model built on the direct coding model elaborated in Lemma 1 with a deterministic analysis transform $T_{A}$ and a deterministic synthesis transform $T_{S}$ (Fig. 1(c)), the following equalities hold automatically:
$I(\bm{X};\bm{\hat{X}})=H(\bm{\hat{X}})$ (8)
$H(\bm{\hat{X}})=H(\bm{U})-H(\bm{U}|\bm{\hat{X}})$ (9)
$H(\bm{U})=I(\bm{X};\bm{\hat{X}})+H(\bm{U}|\bm{\hat{X}})$ (10)
Proof.
Compared to the proof of Lemma 1, for the transform coding, we still have $H(\bm{\hat{X}}|\bm{X})=0$ and $H(\bm{\hat{X}}|\bm{U})=0$. However, since the synthesis transform $T_{S}$ is not guaranteed to be a bijection mapping, the $H(\bm{U}|\bm{\hat{X}})=0$ does not always hold, leading to the correction of Eqn.(9) and thus Eqn.(10). For more details, readers may refer to the Supplementary Material (Sec. 7). ∎
Theorem 2.
For the transform coding model elaborated in Lemma 2, we have
$H(\bm{U})=H(\bm{X})-H(\bm{X}|\bm{\hat{X}})+H(\bm{U}|\bm{\hat{X}}).$ (11)
Proof.
This is immediately apparent from Eqn.(10). ∎
Remarks.
Two conclusions can be drawn for the transform coding model.
  1. (a)
    Equivalence of minimization and maximization. From Theorem 2, for any known source $\bm{X}$, minimizing $H(\bm{U})$ can be equivalently achieved by maximizing the $H(\bm{X}|\bm{\hat{X}})-H(\bm{U}|\bm{\hat{X}})$. For a bijection synthesis function $T_{S}$, of which $H(\bm{U}|\bm{\hat{X}})$ is zero, this maximization objective is reduced to $H(\bm{X}|\bm{\hat{X}})$, which is the same as the direct coding;
  2. (b)
    Achieving $R(D)$ by minimizing $H(\bm{U})$. From Eqn.(10), minimizing $H(\bm{U})$ is simultaneously minimizing $I(\bm{X};\bm{\hat{X}})$ and $H(\bm{U}|\bm{\hat{X}})$. In the ideal case, the informational lower bound given in Eqn.(1) is achieved with $H(\bm{U}|\bm{\hat{X}})$ being reduced to 0.
Figure 2: An illustration of the proposed regularization method, wherein an additional source entropy model $q_{\theta}$ is introduced. The $\mathrm{AE}$ and $\mathrm{AD}$ represent arithmetic encoding and arithmetic decoding, respectively. The side information branch (i.e., the $\bm{\hat{Z}}$ branch) is considered as part of the latent entropy model, which necessitates bits transmission.
Algorithm 1 Training Strategy
1:for each training step do
2:     Sample a minibatch of training images.
3:     Perform the forward pass of the compression network and source entropy model $q_{\theta}$.
4:     Compute the backward gradient and update the compression network based on Eqn.(12).
5:     Freeze the compression network.
6:     Compute the backward gradient and update the source entropy model $q_{\theta}$ based on Eqn.(13).
7:end for
3.2 Information-theoretic regularizer
In principle, optimization neural compression networks with a quantization surrogate, e.g., AUN, introduces gradient estimation errors [14], thereby deviating the objective of minimizing the latent entropy $H(\bm{U})$ under certain distortion constraints (Eqn.(2)), harming both the efficiency and efficacy of the optimization. Motivated by the equivalence of the minimization and maximization optimization discussed in Sec. 3.1, we propose to address this issue by imposing structural maximization constraints on the optimization process. As such, the vanilla objective of minimizing $H(\bm{U})$ is further enhanced through this maximization regularization. Herein, from Theorem 2, the full maximization objective for the transform coding is $H(\bm{X}|\bm{\hat{X}})-H(\bm{U}|\bm{\hat{X}})$, which can be further interpreted as maximizing $H(\bm{X}|\bm{\hat{X}})$ while minimizing $H(\bm{U}|\bm{\hat{X}})$. However, accurately modeling an entropy model $q_{\theta}(\bm{U}|\bm{\hat{X}})$ together with the compression network is quite challenging, as it involves learning another “encoder” function from the reconstruction $\bm{\hat{X}}$ to the latent $\bm{U}$. We experimentally find that a poorly learned $H(\bm{U}|\bm{\hat{X}})$ regularizer can harm the optimization performance, and that the increased training complexity is also burdensome. Consequently, we include only the $H(\bm{X}|\bm{\hat{X}})$ term for regularization, which is an upper bound for the ground-truth $H(\bm{X}|\bm{\hat{X}})-H(\bm{U}|\bm{\hat{X}})$ objective. The entropy modeling is now much easier since the reconstruction $\bm{\hat{X}}$ is closely aligned with the source $\bm{X}$. The resulting training objective is thus given by:
$\mathop{\text{min}}\ R+\lambda D+\alpha\underbrace{\mathbb{E}_{\bm{X}}[\text{log}\ q_{\theta}(\bm{X}|\bm{\hat{X}})]}_{\approx-H(\bm{X}|\bm{\hat{X}})},$ (12)
where $R+\lambda D$ follows the rate-distortion loss in Eqn.(2); $-H(\bm{X}|\bm{\hat{X}})$ is the proposed regularizer; $q_{\theta}(\bm{X}|\bm{\hat{X}})$ is the estimated source entropy model; $\alpha$ is regularization factor, controlling the degree of regularization. An illustration for this regularization method is given in Fig.(2). In addition to the compression network, $q_{\theta}(\bm{X}|\bm{\hat{X}})$ is introduced, which is parameterized by a neural network and can be jointly trained via the maximum likelihood criteria [41]:
$\mathop{\text{max}}\ \mathbb{E}_{\bm{X}}[\text{log}\ q_{\theta}(\bm{X}|\bm{\hat{X}})].$ (13)
Overall, the joint training strategy for the compression network and $q_{\theta}(\bm{X}|\bm{\hat{X}})$ is summarized in Algorithm 1, wherein a GAN-style two-stage training strategy [42] is employed.
Figure 3: Performance of the proposed regularization method on three compression models, i.e., hyperprior [4], autoregressive [8], and attention [10]. The anchor is training with the vanilla rate-distortion loss (Eqn.(2)) under an equal number of training steps. $\alpha$ indicates the regularization factor (Eqn.(12)).
4 Experiments
4.1 Experimental setup
Compression models.
Three image compression baseline models are used in the experiments: the hyperprior context model [4] with scale and mean [8], the joint hyperprior and spatial autoregressive context model [8], and the attention architecture [10]. We first train these three models from scratch using the vanilla rate-distortion loss (Eqn.(2)). Then, we compare the results of training these models with the proposed regularization method (Eqn.12) under an equal number of training steps. All three models follow CompressAI’s reimplementation [43], wherein the quantization surrogate is ANU. We may refer to the above three models as hyperprior, autoregressive, and attention, respectively.
Conditional source entropy modeling.
For each compression model, we apply exactly the same entropy modeling method as the latent entropy modeling. For instance, for the autoregressive, the $q_{\theta}(\bm{X}|\bm{\hat{X}})$ is also modeled as the Gaussian distribution conditioned on $\bm{\hat{X}}$ with spatial autoregressive context over $\bm{X}$. Our experiments indicate that aligning the entropy model between the latent and source is critical for effective regularization; deviations in either direction (stronger or weaker source models) degrade compression performance. For more details on the network structure of $q_{\theta}(\bm{X}|\bm{\hat{X}})$ and the alignment, readers may refer to the Supplementary Material (Sec. 9 and Sec. 10, respectively).
Training.
For training set, we use Flickr20k [44], consisting of 20745 natural images from the Flickr.com. The train-to-validation ratio is set to 9:1. To remove the effect of randomness and facilitate a fair performance comparison, the random seed is fixed to 1 for all training. Following the settings of CompressAI, we train four bit-rate points for each model, i.e., $\lambda\in\{0.0018,0.0035,0.0067,0.0130\}$. The batch size is 16, and the image patch size is $256\times 256$. For optimizer, the experiments use Adam [45], and the learning rates for the compression model and the extra source entropy model are set to $10^{-4}$ and $10^{-3}$, respectively. Herein, following the performance evaluation protocol in prior works [13, 12, 30], we train each model for $2\times 10^{6}$ steps and report the performance, wherein the original attention paper [10] only trains for $1\times 10^{6}$ steps. The performance comparison between the reproduced baselines and CompressAI’s pre-trained models is elaborated in Supplementary Material (Sec. 8). In summary, compared to the pre-trained, our reproduction models yield 0.39%, 0.41%, and -4.11% BD-Rates [46] for the hyperprior, autoregressive, and attention models, respectively, serving as decent reproduction for the experiments. All experiments are conducted on an Nvidia RTX 4090 GPU.
Performance evaluation.
We follow the convention of evaluating over natural image datasets of Kodak [47], CLIC 2024 Validation Set [48] and Tecnick [49]. Besides, to demonstrate generalization performance, we evaluate the models on four out-of-domain datasets: (a) pixel-style dataset [50] (100 images in total); (b) screen-content dataset SCI1K [51] (the first 100 images of the test set); (c) game dataset CCT-CGI [52] (24 images in total); (d) pathology dataset BRACS [53] (the first 100 images of the test set).
4.2 Results
Performance on natural images.
In Fig.3, we present the BD-Rate comparison at four training steps: $0.5\times 10^{6}$, $1\times 10^{6}$, $1.5\times 10^{6}$, and $2\times 10^{6}$. The regularization factors $\alpha$ for the hyperprior, autoregressive, and attention models are set to 0.1, 1 and 1, respectively. It can be observed that, at most equal training steps for all three models, our proposed regularization method can yield better compression performance than the vanilla optimization method. In particular, at the $2\times 10^{6}$ step, the BD-Rates for hyperprior, autoregressive, and attention are -0.88%, -0.92% and -0.92%, respectively. For hyperprior and autoregressive, after a brief convergence period, the BD-Rates are gradually improved from 0 to around -1%. For attention, the regularization immediately improves the compression performance at a very early $0.5\times 10^{6}$ training step. In Fig.4, we present the detailed rate-distortion curves at the $2\times 10^{6}$ step on the Kodak dataset. It can be observed that, due to the introduction of the additional conditional source entropy regularization, the bit rate (or the latent entropy) of each point is decreased by a smaller value. From the form of regularization in Eqn.(12), i.e., $-\alpha H(\bm{X}|\bm{\hat{X}})$, the extent of such bit rate reduction depends on both the regularization factor $\alpha$ and the conditional source entropy $H(\bm{X}|\bm{\hat{X}})$. Note that a more advanced conditional source entropy modeling for $q_{\theta}(\bm{X}|\bm{\hat{X}})$ does not experimentally improve the optimization performance. Instead, the alignment with the latent entropy model proves to be more impactful.
Figure 4: Compression performance on Kodak. Our regularizer achieves -1.13%, -1.15%, and -0.84% BD-Rates for the hyperprior, autoregressive, and attention models, respectively.
Table 1: Generalization performance of the proposed regularization method on four out-of-domain datasets, i.e., the pixel-style [50], screen [51], game [52], and pathology [53] datasets. The in-domain performance on natural images [47, 48, 49] is also listed for reference.
Test
set
Regularized v.s. Unregularized
hyperprior autoregressive attention
Pixel-style -2.20% -1.96% -0.47%
Screen -1.15% -2.01% -0.69%
Game -1.31% -0.83% -0.82%
Pathology -1.11% -0.81% -2.38%
Natural -0.88% -0.92% -0.92%
Generalizing on unseen domains.
In machine learning, regularization plays an essential role in preventing overfitting and enhancing generalization. In Table 1, we directly apply the models trained on natural images (at the $2\times 10^{6}$ step) to the four unseen domains and evaluate the effectiveness of our regularization method. The results indicate that, across all new domains, regularized models consistently outperform unregularized ones. Performance improvements vary by domain and model, with notable gains in some cases. For instance, in the pathology domain with the attention model, the regularization achieves a BD-Rate of -2.38%, more than double the improvement observed in the natural domain.
More experiments.
In Fig.(5), we continue training both the unregularized and regularized hyperprior models to $4\times 10^{6}$ steps and demonstrate the effect of the regularization factor $\alpha$ on the regularization. Herein, with extended training, the regularization consistently enhances optimization. The best performance, with $\alpha=1$, is achieved at the $2.5\times 10^{6}$ step, showing a BD-Rate of -1.24%. As training progresses to $4\times 10^{6}$ steps, the gain reduces to around 0.7%. For the effect of hyperparameter $\alpha$, it is shown that either the value is too large or too small will not benefit the optimization, which can be explained as follows. As discussed in Sec. 3.2, our $\alpha H(\bm{X}|\bm{\hat{X}})$ maximization regularization is essentially a bound upper of the ground-truth objective $\alpha(H(\bm{X}|\bm{\hat{X}})-H(\bm{U}|\bm{\hat{X}}))$. When $\alpha$ is large, this $H(\bm{U}|\bm{\hat{X}})$ gap is also amplified, which harms the regularization accuracy and thus the optimization performance. When $\alpha$ is small, the regularizer has little contribution to the overall loss, and the optimization is reduced to the vanilla objective.
Figure 5: Effects of regularization factor $\alpha$. The compression model is hyperprior and the anchor is training with the vanilla rate-distortion loss (Eqn.(2)) under an equal number of training steps.
Complexity.
Our regularization method is only applied at the network training stage and thus imposes no added inference complexity. As for training, the main overhead of our method comes from training an extra source entropy model $q_{\theta}(\bm{X}|\bm{\hat{X}})$, whose network structure is intentionally aligned with the latent entropy model of CompressAI’s reimplementation in our experiment. For the case of $\lambda=0.0018$, we collect the absolute training time of each model and calculate the ratio. As shown in Table. 2, compared to the unregularized anchor, the training time of our proposed method is increased by 19.5%, 11.3%, and 1.5% for hyperprior, autoregressive, and attention, respectively. Notably, as the compression backbone becomes more complex, the overhead is less prominent.
Table 2: Training complexity of the proposed regularizer compared to the unregularized anchor.
hyperprior autoregressive attention
Training time 119.5% 111.3% 101.5%
5 Conclusion
This paper investigates a structural entropy regularizer for lossy neural image compression. Specifically, motivated by the information-theoretic analysis of compression models, we demonstrate that, in addition to minimizing latent entropy, maximizing conditional source entropy can be equally important. Subsequently, by regularizing the neural networks to simultaneously maximize the conditional source entropy during end-to-end optimization, we experimentally observe improvements in both the in-domain compact representation and the out-of-domain generalization abilities of compression models. Additionally, as the primary overhead of this regularizer comes from training an additional source entropy model, the proposed method imposes no added inference complexity, and the training complexity remains negligible when the compression backbone is substantially more complex than the source entropy model, opening up a new space of regularizing neural compression networks.
Supplementary Material
6 Proof of Lemma 1
For a deterministic quantization process ${\rm Q}(\cdot)$, the conditional probability $p(\bm{U}|\bm{X})$ can only take values of 0 or 1, i.e.,
$p(\bm{U}|\bm{X})=\left\{\begin{array}[]{lr}1,\quad\text{if}\ \ \bm{U}={\rm Q}(\bm{X})\\ 0,\quad\text{if}\ \ \bm{U}\neq{\rm Q}(\bm{X})\end{array}\right..$ (14)
Similarly, for a deterministic dequantization process ${\rm Q}^{-1}(\cdot)$, it is by definition a bijection function, i.e., given the index $\bm{U}$ or the reconstruction $\hat{\bm{X}}$, we can uniquely determine the corresponding value of $\hat{\bm{X}}$ or $\bm{U}$, respectively. Therefore, we have
$p(\hat{\bm{X}}|\bm{U})=\left\{\begin{array}[]{lr}1,\quad\text{if}\ \ \hat{\bm{X}}={\rm Q}^{-1}(\bm{U})\\ 0,\quad\text{if}\ \ \hat{\bm{X}}\neq{\rm Q}^{-1}(\bm{U})\end{array}\right.,$ (15)
$p(\bm{U}|\hat{\bm{X}})=\left\{\begin{array}[]{lr}1,\quad\text{if}\ \ \bm{U}={\rm Q}(\hat{\bm{X}})\\ 0,\quad\text{if}\ \ \bm{U}\neq{\rm Q}(\hat{\bm{X}})\end{array}\right.,$ (16)
and overall,
$p(\hat{\bm{X}}|\bm{X})=\left\{\begin{array}[]{lr}1,\quad\hat{\bm{X}}={\rm Q}^{-1}({\rm Q}(\bm{X}))\\ 0,\quad\hat{\bm{X}}\neq{\rm Q}^{-1}({\rm Q}(\bm{X}))\end{array}\right..$ (17)
Thus, given Eqn.(15) to Eqn.(17), the following conditional entropy is by definition equal to 0:
$H(\bm{\hat{X}}|\bm{U})=H(\bm{U}|\bm{\hat{X}})=H(\bm{\hat{X}}|\bm{X})=0.$ (18)
Furthermore, from the equality of mutual information $I(\bm{X};\bm{\hat{X}})$ and the chain rule of joint entropy $H(\bm{\hat{X}},\bm{U})$, we have
$\left\{\begin{array}[]{lr}I(\bm{X};\bm{\hat{X}})=H(\bm{\hat{X}})-H(\bm{\hat{X}}|\bm{X})\\ H(\bm{U}|\bm{\hat{X}})+H(\bm{\hat{X}})=H(\bm{\hat{X}}|\bm{U})+H(\bm{U})\end{array}\right..$ (19)
Substituting Eqn.(18) into Eqn.(19), we can conclude
$\left\{\begin{array}[]{lr}I(\bm{X};\bm{\hat{X}})=H(\bm{\hat{X}})\\ H(\bm{\hat{X}})=H(\bm{U})\end{array}\right.,$ (20)
and thus
$H(\bm{U})=I(\bm{X};\bm{\hat{X}}).$ (21)
7 Proof of Lemma 2
Recalling the proof in Sec. 6, since both the analysis transform $\mathrm{T}_{\mathrm{A}}(\cdot)$ and synthesis transform $\mathrm{T}_{\mathrm{S}}(\cdot)$ are deterministic, we still have
$p(\bm{U}|\bm{X})=\left\{\begin{array}[]{lr}1,\quad\text{if}\ \ \bm{U}={\rm Q}(\mathrm{T}_{\mathrm{A}}(\bm{X}))\\ 0,\quad\text{if}\ \ \bm{U}\neq{\rm Q}(\mathrm{T}_{\mathrm{A}}(\bm{X}))\end{array}\right.,$ (22)
$p(\hat{\bm{X}}|\bm{U})=\left\{\begin{array}[]{lr}1,\quad\text{if}\ \ \hat{\bm{X}}=\mathrm{T}_{\mathrm{S}}({\rm Q}^{-1}(\bm{U}))\\ 0,\quad\text{if}\ \ \hat{\bm{X}}\neq\mathrm{T}_{\mathrm{S}}({\rm Q}^{-1}(\bm{U}))\end{array}\right.,$ (23)
$p(\hat{\bm{X}}|\bm{X})=\left\{\begin{array}[]{lr}1,\quad\text{if}\ \ \hat{\bm{X}}=\mathrm{T}_{\mathrm{S}}({\rm Q}^{-1}({\rm Q}(\mathrm{T}_{\mathrm{A}}(\bm{X}))))\\ 0,\quad\text{if}\ \ \hat{\bm{X}}\neq\mathrm{T}_{\mathrm{S}}({\rm Q}^{-1}({\rm Q}(\mathrm{T}_{\mathrm{A}}(\bm{X}))))\end{array}\right.,$ (24)
and therefore
$H(\bm{\hat{X}}|\bm{U})=H(\bm{\hat{X}}|\bm{X})=0.$ (25)
Herein, the core distinction between the transform coding and direct coding models lies in the fact that the synthesis transform $\mathrm{T}_{\mathrm{S}}(\cdot)$, unlike the dequantization function ${\rm Q}^{-1}(\cdot)$, does not inherently guarantee a bijective mapping, particularly in the context of neural transforms. Consequently, given the reconstruction $\hat{\bm{X}}$, there may be uncertainty in $\bm{\hat{Y}}$ and thus $\bm{U}$, i.e.,
$H(\bm{U}|\bm{\hat{X}})\neq 0.$ (26)
Substituting Eqn.(25) and Eqn.(26) into Eqn.(19), we can conclude
$\left\{\begin{array}[]{lr}I(\bm{X};\bm{\hat{X}})=H(\bm{\hat{X}})\\ H(\bm{\hat{X}})=H(\bm{U})-H(\bm{U}|\bm{\hat{X}})\end{array}\right.,$ (27)
and thus
$H(\bm{U})=I(\bm{X};\bm{\hat{X}})+H(\bm{U}|\bm{\hat{X}}).$ (28)
8 Reproduced baselines
Table 3: BD-Rate comparison between our reproduction and the pre-trained models.
Kodak Tecnick CLIC Average
hyperprior 0.52% 1.88% -1.24% 0.39%
autoregressive 0.56% 2.25% -1.56% 0.41%
attention -3.32% -3.33% -5.68% -4.11%
We retrain the hyperprior [4], autoregressive [8] and attention [10] models from scratch, adhering to the default implementation and training configurations of CompressAI’s [43]. Four bit-rate points, i.e., $\lambda\in\{0.0018,0.0035,0.0067,0.0130\}$ are trained with $2\times 10^{6}$ steps. The evaluation results are summarized in Table 3. On average, when compared to the pre-trained models of CompressAI’s, our reproduction models yield 0.39%, 0.41%, and -4.11% BD-Rates [46] for the hyperprior, autoregressive, and attention models, respectively. Notably, on the CLIC dataset, our reproduction models outperform their pre-trained counterparts by -1.24%, -1.56%, and -5.68% BD-Rates for the hyperprior, autoregressive, and attention models, respectively. For the attention model, our reproduction model outperforms the pre-trained model on all three evaluation datasets of Kodak, CLIC, and Tecnick by -3.32%, -3.33%, and -5.68% BD-Rates, respectively. The detailed rate-distortion curves for the Kodak dataset are visualized in Fig.6.
Figure 6: Comparison between the pre-trained models from CompressAI’s and our reproduction on the Kodak dataset. Our reproduction achieves 0.52%, 0.56%, and -3.32% BD-Rates for the hyperprior, autoregressive, and attention models, respectively.
9 Conditional source entropy models
Similar to the latent entropy modeling [4, 8, 10], our regularizer involves an additional modeling of the conditional distribution $q_{\theta}(\bm{X}|\bm{\hat{X}})$. Herein, $q_{\theta}(\bm{X}|\bm{\hat{X}})$ follows exactly the same modeling method and the neural network structure as the latent entropy used. In particular, for the hyperprior model, we have
$q_{\theta}(\bm{X}|\bm{\hat{X}})\sim\prod\limits_{i=1}^{N}\mathbb{N}(\mu_{i},\sigma_{i}^{2}),$ (29)
$\text{where}\ (\mu_{1},...,\mu_{N},\sigma_{1}^{2},...,\sigma_{N}^{2})=f_{\theta}(\bm{\hat{X}}),$ (30)
and $f_{\theta}$ is the neural network; $\mathbb{N}$ denotes the Gaussian distribution; $\mu_{i}$ and $\sigma_{i}^{2}$ denote the mean and variance for the $i^{th}$ dimension, respectively; $N$ denotes the dimension of both $\bm{X}$ and $\bm{\hat{X}}$. For both the autoregressive and attention models, the distribution remains Gaussian. However, for the mean and variance, they are determined by the extra spatial context, i.e.,
$(\mu_{1},...,\mu_{N},\sigma_{1}^{2},...,\sigma_{N}^{2})=f_{\theta}(\bm{\hat{X}},\bm{X}_{<i}),$ (31)
where $\bm{X}_{<i}$ denotes the pixels before the current position. Note that in the original attention network [10], the latent is modeled as the mixture of Gaussian with 3 clusters. In CompressAI’s default implementation, this cluster number is simplified to 1. We follow this latent implementation and use an aligned 1-cluster conditional Gaussian distribution for the source modeling. The detailed network structure for $q_{\theta}(\bm{X}|\bm{\hat{X}})$ is depicted in Fig.7, wherein the spatial context is achieved through a $5\times 5$ masked convolution layer (i.e., MaskConv).
Figure 7: Conditional source entropy models for (a) hyperprior; (b) autoregressive and attention; (c) Details of Conv Block 1; (c) Details of Conv Block 2.
10 On the alignment of entropy models
Figure 8: Regularization performance for (a) the hyperprior compression network with the autoregressive regularizer; (b) the autoregressive compression network with the hyperprior regularizer.
In Sec.4, we highlight the benefits of aligned entropy modeling between latent and source distributions. In this section, we extend the analysis to explore cases where the entropy models are unaligned, investigating two scenarios: a compression network equipped with (a) a stronger source entropy model and (b) a weaker source entropy model. In particular, for the entropy models of the hyperprior and autoregressive networks discussed in Sec.9, we consider the autoregressive entropy model with additional spatial contexts to be a more accurate and thus “stronger” model. Thereafter, using the same experimental setting in Sec.4, we train the hyperprior compression network with a autoregressive source regularizer and the autoregressive compression network with a hyperprior source regularizer, and the results are depicted in Fig.8(a) and Fig.8(b), respectively. The values of $\alpha$ are still set to 0.1 and 1 for the hyperprior and autoregressive regularizers, respectively. The performance of the aligned entropy model is also included for reference. From Fig.8, it can be observed that neither a stronger nor a weaker source entropy model consistently facilitates the optimization performance of the compression networks. Specifically, when a stronger autoregressive regularizer is imposed on the hyperprior network, the average BD-Rate is always above zero during the training. When a weaker hyperprior regularizer is employed by the autoregressive network, the optimization is temporarily improved at the $1.5\times 10^{6}$ training step, with a BD-Rate of -0.20%. In contrast, by simply aligning the entropy models, significant performance improvements are observed.